Please Help
A forester mixes gasoline and oil to make 2 gallons of mixture for his two-cycle chain-saw engine. This mixture is 32 parts gas and 1 part two-cycle oil. How much gas must be added to bring the mix to 40 parts gas and 1 part oil.
Thanks

Answer by stanbon(48558)

Let number of gallons of pure gas be "x".
Follow the gas:
Start with (32/33)(2 gallons)
Add x gallons
End up with (40/41)(2+x)gallons
EQUATION:
(32/33)(2)+x=(40/41)(2+x)
Multiply through by 33*41 to get:
2624 +1353x=2640+1320x
33x=16
x=16/33=0.4848 gallons
Cheers,
Stan H.